You have a liquid culture of E. coli at 7.5 x 10^9 org/mL concentration. Draw a serial dilution scheme to give a final concentration of 1.5 x 10^3 org/mL and indicate the volume of sample, volume of diluent, DF, TDF, and the concentration of each tube. Please explain steps. I am not too sure how to even start problems like these. Thank you!
Answer #1

7.5 X109    org/ml

Take 1ml add 4 ml of medium ( so I ml of culture in 5 ml of volume , 1:5 dilution)

Dilution factor =final volume / volume of the culture = 5/1=5

7.5 X 109 org/ml   is diluted 5 times so the new concentration = (7.5 X 109)/5 = 1.5 X109 org/ml

Take 0.1 ml (100 µl) from this and make up the volume to 10 ml (add 9.9 ml of medium) (1:100 dilution)

Dilution factor =final volume / volume of the culture = 10/0.1=100

1.5 X109 org/ml   is diluted 100 times so the new concentration = (1.5 X 109)/100 = 1.5 X107 org/ml

Take 0.1 ml (100 µl) from this and make up the volume to 10 ml (add 9.9 ml of medium) (1:100 dilution)

Dilution factor =final volume / volume of the culture = 10/0.1=100

1.5 X107 org/ml   is diluted 100 times so the new concentration = (1.5 X 107)/100 = 1.5 X105 org/ml

Take 0.1 ml (100 µl) from this and make up the volume to 10 ml (add 9.9 ml of medium) (1:100 dilution)

Dilution factor =final volume / volume of the culture = 10/0.1=100

1.5 X105 org/ml   is diluted 100 times so the new concentration = (1.5 X 105)/100 = 1.5 X103 org/ml